y=2x+16 2x-5y=-16 need to find what is x and y
Answer by mananth(16949)

y=2x+16
add -2x to both sides
y-2x=2x-2x+16
y-2x=16 .........1
2x-5y=-16........2
..
multiply equation 1 by 5 and add equation 2 to it
5y-10x=80
2x-5y=-16
add them
2x-10x=80+16
-8x=96
divide by -8
-8x/-8=96/-8
x=-12
plug the value of x in the equation 1
y-(-12)(2)=16
y+24=16
add -24 on both sides
y+24-24=16-24
y=-8